If A, B are differentiable vector point function of scalar variable f over domain S, then prove that d/dt( AxB) = (dA/dt x B) + (A x dB/dt) 


1
Expert's answer
2021-03-31T01:55:26-0400

Given A, B are differentiable vector point functions of scalar variable f over domain S.

then we have to prove that: "\frac{d(A \times B)}{dt}" = "\frac{dA}{dt} \times B" + "A \times \frac{dB}{dt}"

Let A = "<" x1, y1, z1 ">" and B = "<" x2, y2, z2 ">"

"\implies" "\frac{dA}{dt}" = "<" x1', y1', z1' ">" and "\frac{dB}{dt}" = "<" x2', y2', z2' ">"

We have,

A "\times" B = "\begin{vmatrix}
 i & j & k \\
 x_1 & y_1 & z_1\\
 x_2 & y_2 & z_2
 
\end{vmatrix}" = = "<" y1z2 − z1y2, z1x2 − x1z2, x1y2 − y1x2 ">"


Taking derivatives using the product rule from single variable calculus, we get a lot of terms, which we can group to prove the vector formula.


"\frac{d(A \times B)}{dt}" = "<" y1'z2 + y1z2' − z1'y2 − z1y2', z1'x2 + z1x2' − x1'z2 − x1z2', x1'y2 + x1 y2' − y1'x2 − y1x2' ">"

= "<" (y1'z2 − z1'y2)+ (y1z2'− z1y2'), (z1'x2 − x1'z2) + (z1x2'− x1z2'), (x1'y2 − y1'x2) + (x1y2'− y1x2') ">"

= "<" x1', y1', z1' ">" "\times" "<" x2, y2, z2 ">" + "<" x1, y1, z1 ">" "\times" "<" x2', y2', z2' ">"


= "\frac{dA}{dt} \times B" + "A \times \frac{dB}{dt}"

Hence proved,


"\frac{d(A \times B)}{dt}" = "\frac{dA}{dt} \times B" + "A \times \frac{dB}{dt}"